# HADIS: Hybrid Adaptive Diffusion Model Serving for Efficient Text-to-Image Generation

### v1

@@ 1. Introduction
Model Serving Goals. A model serving system has two key goals. First, the system must use resources efficiently to serve queries (i.e., text prompts from the user) with high throughput. Second, since queries are served to users in real time, the system must guarantee service-level objectives (SLOs) that require the latency of response to be within a specified threshold, even as user queries vary widely in complexity. Simple queries like “a banana on a white plate” require minimal computation, whereas complex ones like “a surreal cityscape representing time and memory” demand deeper semantic understanding and finer generative detail. Serving all queries with a “heavyweight” model that has a high computational cost leads to excessive resource usage and response latencies, while using a low-cost “lightweight” model alone may compromise the quality of the response. Therefore, a model serving system must dynamically adjust to query arrival rates and query difficulty to satisfy overall throughput and SLO constraints.
@@
Model Scaling. Model scaling has emerged as a key approach to designing model serving systems. An initial example of this approach proposed in Proteus (Ahmad et al., 2024) introduces the notion of accuracy scaling that uses less-accurate lower-cost “lightweight” models during peak periods when higher throughput is required, while using more accurate higher-cost models at other times. More recently, query awareness has been used to enhance model scaling where each query is processed through a model cascade that is a set of models sequenced from lighter weight ones to heavier weight ones. Each query is routed sequentially through the models in the cascade till a response of acceptable quality is obtained. DiffServe (Ahmad et al., 2025) is an example of such a cascade using two stages: every prompt is first processed by a lightweight model, and the generated image is assessed by a discriminator. If the image is deemed low-quality, the prompt is re-routed to a heavyweight model. The easier queries only require the first stage of the cascades, saving resources and latency, while the harder queries would need to be routed to the second.
@@
Limitations of Current Solutions. The current state-of-the-art serving systems using model cascades have two key limitations. First, the models chosen for a cascade may not be optimal for all query workloads. For instance, DiffServe relies on a fixed pair of lightweight and heavyweight models, even though different workloads and latency budgets may favor different combinations. Second, model cascades inherently increase both resource usage and latency as some fraction of the queries need to get routed through multiple models. For instance, in a two-model cascade, all prompts are sent to a lightweight model, which wastes resources for queries that are hard and require heavyweight processing regardless.
@@
(i) Dynamic choice of models. As diffusion models proliferate, each variant offers different trade-offs between latency and quality. Statically fixing model combinations in a cascade across all query workload scenarios is inefficient. Therefore, we design a model serving system that dynamically chooses the models in a cascade to suit the properties of the current query workload.
@@
(ii) Resource-efficient query routing. Routing each query through the entire model cascade wastes computational resources and increases the latency. For instance, routing the harder queries to a lightweight model is unlikely to produce a response of acceptable quality, leading to a waste of compute resources and an increase in the latency. We propose a hybrid architecture shown in Figure 1 that consists of both a router and a discriminator to reduce resource wastage and latency. Our system begins with a rule-based router that computes a query difficulty score using linguistic and semantic features, including token rarity, abstractness, spatial relations, etc. Queries that are predicted to be hard are routed directly to a heavyweight model, skipping the lightweight model(s) in the cascade. Queries that are classified as easy are processed by a lightweight model, whose outputs are then evaluated by an image discriminator for both semantic and aesthetic quality. If the generated images fail to meet the quality requirements, the query is sent to the heavyweight model. This hybrid design reduces computational overhead and latency for easy queries and avoids wasteful lightweight model inference on hard queries.
@@
(iii) Managing cascade complexity. The plethora of diffusion model variants, each offering a different quality-latency tradeoff, raises the possibility of complex model serving architectures consisting of many model variants, discriminators and router components. We show that such complexity is unnecessary and it is sufficient to dynamically choose a simple two-stage model cascade with a single router and discriminator to achieve the benefits of model scaling. This observation greatly reduces the space of architectural choices and enables efficient resource management.
@@
(iv) Reducing the complexity of resource management. Resource management of model cascades is a complex NP-hard problem that is often solved using a Mixed-Integer Linear Program (MILP) (Ahmad et al., 2025). While restricting the architecture to two-stage cascades helps, we reduce the combinatorial complexity even further as follows. We use offline profiling to benchmark all pairwise model combinations under different routing thresholds and generate a cascade configuration lookup table. Each configuration in this table encodes a specific combination of models, thresholds, latency measurements, and the estimated ratio of hard queries (i.e., the percentage of queries that will likely be routed to a heavyweight model given the current threshold). The table prunes the search space by identifying Pareto-optimal configurations that offer the best quality-latency trade-offs. At runtime, the system leverages this table to select a configuration that meets the desired SLO constraints. To determine the optimal deployment strategy given a workload distribution and system constraints (e.g., GPU count, latency budget, etc), we formulate the configuration selection and resource allocation as an MILP problem. This formulation jointly optimizes routing decisions, model assignments, and resource partitioning to maximize overall serving efficiency and quality.
Figure 1. Text-to-Image Serving Workflow of HADIS.
@@
  • We propose adaptive model cascade selection, dynamically selecting optimal pairwise model cascades based on various workload and latency constraints.
  • We propose a hybrid routing strategy consisting of a rule-based query router and a learned image discriminator, jointly reducing computational overhead while preserving response quality.
  • We address the complex resource management problem for adaptive model cascade through profiling Pareto-optimal configurations and formalizing it as a mixed integer linear programming (MILP) problem.
@@
2. Background and Rationale for Design
We discuss the background and rationale for the key design choices made by our model serving system, HADIS.
2.1. Background for text-to-image diffusion models
@@
Text-to-image diffusion models embody a clear quality-latency trade-off (see Figure 2-left): more compute (e.g., deeper networks or transformers, more denoising steps, etc) typically yields better image quality but longer runtime (and latency) and lesser throughput (i.e., fewer queries processed per second). In our setting on L40S GPUs, the spectrum spans sub-second generation (e.g., SDXL-Lightning at 2 steps $\approx$ 0.5s) to tens of seconds (e.g., SD3.5-Large at 50 steps $\approx$ 27s) for an image of $\times1024$ resolution.
@@
Figure 2. Quality-latency tradeoffs of different diffusion model variants (left) and model cascades (right). Quality is measured by Fréchet Inception Distance (Heusel et al., 2017) (FID, lower is better). The cascades are constructed by the model variants colored in orange in the left panel. The orange dashed lines partition latency into three regimes where each regime has a distinct model cascade that yields the frontier points.
Prior query-aware model serving approaches (Ahmad et al., 2025; Chen et al., 2023) route queries via a cascade to optimize throughput and SLO violation ratios. In the case of a simple two-model cascade, the first model could be a lightweight model that offers low latency and relatively low quality, and the second model could have higher latency and quality. A discriminator decides if the output of the first model is of sufficiently high quality, failing which the query is routed to the second model of the cascade. This reduces average latency relative to always using a heavyweight model, but still incurs a non-trivial overhead since every query incurs the overhead of the first lightweight model. Further, a longer cascade with more than two models could incur more computational and latency overheads as some queries would pass through all models in the cascade to produce a response.
2.2. Rationale for two-model cascades
@@
We show that two-model cascades offer comparable quality-latency trade-offs as model cascades with more model variants. Figure 2-right explores the design space of diffusion-model cascades using four model variants (orange stars in the left panel). For each candidate cascade we enumerate routing policies specified by the fraction of queries routed to the heavyweight model(s), and measure the resulting mean latency (x-axis) and FID (y-axis) on 5k prompts. The right panel plots all feasible cascade configurations in grey by varying model combinations and routing policies. The Pareto frontier [^1] A Pareto frontier $P$ is a set of model cascades such that for every $m\in P$ there is no other model cascade $m^{\prime}\in P$ that has both lower FID and latency than $m$. In that sense, all model cascades in the Pareto frontier represent an optimal tradeoff between the two objectives where one objective can be made better only at the cost of the other. obtained by two-model and three-model cascades are colored in red and blue, respectively. Two empirical facts stand out. First, the red and blue frontiers coincide over the relevant latency range: for any latency budget, a two-model cascade attains essentially the same best FID as a three-model cascade. Second, while the blue frontier can fill in a few intermediate points (i.e., slightly finer granularity), it does not deliver significantly better quality at a given latency.
@@
Figure 3. Limitations of fixed model cascades. In both panels, each point represents a distinct two-model cascade. Comparing between low load (left) and high load (right), the best cascade (circled in red) varies as the system workload and the tolerable SLO violation ratio change.
2.3. Rationale for dynamic model choice
Given that our system adopts a two-model cascade, it is natural to ask which pair of models can provide the best system performance. We empirically observe that none of the fixed lightweight/heavyweight model pairs remains optimal across varying query demands.
Figure 2-right shows the best two-model cascade configurations vary with the latency budget. It plots the Pareto frontier of two-model cascade configurations built from four candidates in the left panel. The orange dashed lines partition latency into three regimes. Within each regime, the frontier points arise from a different model cascade, ¡SDXL-Lightning, SD3.5-Turbo¿ at low latency, ¡SD3.5-Turbo, SD3.5-Medium¿ at mid latency, and ¡SD3.5-Medium, SD3.5-Large¿ at high latency. We observe that different model cascades offer different latencies and response qualities. Hence, when the allowed SLO is tighter, a faster cascade is preferred to satisfy deadlines; when the SLO relaxes, cascades of higher quality can be chosen to increase the response quality.
Figure 3 further shows the best two-model cascades vary with the query loads. It compares several two-model cascades under low demand (left panel) and high demand (right panel) with the SLO held fixed. Each cascade serves queries with the resource allocation algorithm in DiffServe (Ahmad et al., 2025). At low load, all pairs keep SLO violations small. ¡SD3.5-Turbo, SD3.5-Medium¿ (red star) attains the best violation rate with competitive FID, while ¡SD3.5-Medium, SD3.5-Large¿ trades a slight increase in violations for the best FID (grey diamond). Under high load, as capacity constraints surface, pairs involving slower models (i.e., grey diamond) cannot meet throughput demands and incur large violations, whereas ¡SDXL-Lightning, SD3.5-Turbo¿ (blue down-pointing triangle) remains within SLO and delivers the best quality–latency balance for that demand.
@@ 2.4. Rationale for hybrid query routing architecture
We argue for a hybrid query routing architecture by noting that both a router and a discriminator are needed and either one alone is insufficient, as they address complementary failure modes of the model cascade. If we want to solely use a router to evaluate the hardness of queries and route queries based on that, the router should be lightweight, introducing negligible overhead, and also accurate such that most of the queries will not be misclassified, i.e., false “hard” queries can cause unnecessary waste of GPU resources, and false ”easy” queries can reduce the response quality. While query hardness has been studied in  (Xing et al., 2010; De Rosa Palmini and Cetinic, 2025; Liu and Chilton, 2022), estimating query hardness for text-to-image generation with a router is still an open and non-trivial research problem. Building a router that classifies queries with high accuracy is difficult.
@@
A discriminator in the model cascade is usually effective in determining the hardness of a query, which is done in prior work (Ahmad et al., 2025; Chen et al., 2023). However, if we use a discriminator alone, all queries will go through the lightweight model regardless of hardness, then be evaluated by the discriminator, which brings in additional overhead, especially when the first model of the cascade requires significant computation. For instance, in Figure 2, when adopting model pair ¡SD3.5-Medium, SD3.5-Large¿, of which the latency is $\sim$13s and $\sim$27s respectively, the additional overhead for the first model of the cascade can be 33% of the execution latency.
@@
We therefore adopt a hybrid architecture, combining both a router and a discriminator. The rule-based router sends obviously hard queries directly to the heavyweight model, while the rest go to the lightweight model. The discriminator then evaluates the outputs from the lightweight model and catches false “easy” cases for re-routing to heavyweight models. This design avoids unnecessary first-stage work for clearly hard queries, preserves quality by correcting router mistakes, and keeps the added control cost minimal. Our evaluation in Section 4.4 shows this hybrid routing reduces SLO violations versus discriminator-only and improves FID versus router-only, delivering the best quality-latency trade-off under identical hardware resources.
3. HADIS Design

### v2

@@ 1 Introduction
However, the computational cost of these models remains a significant bottleneck for real-time and scalable deployments in production model serving systems. Generating a single image often requires multiple seconds of GPU time due to the iterative denoising process. While recent few-step diffusion variants such as SDXL-Lightning [26] and SDXL-Turbo [37] have been explicitly designed for real-time or near-real-time generation on commodity hardware, this work addresses the problem from the model serving perspective.
@@
Model Serving Goals. A model serving system has two key goals. First, the system must use resources efficiently to serve queries (i.e., text prompts from the user) with high throughput. Second, since queries are served to users in real time, the system must guarantee service-level objectives (SLOs) that require the latency of response to be within a specified threshold, even as user queries vary widely in complexity. Simple queries like “a banana on a white plate” require minimal computation, whereas complex ones like “a surreal cityscape representing time and memory” demand deeper semantic understanding and finer generative detail. Serving all queries with a “heavyweight” model that has a high computational cost leads to excessive resource usage and response latencies, while using a low-cost “lightweight” model alone compromises the response quality. Therefore, a model serving system must dynamically adjust to query arrival rates and query difficulty to satisfy both throughput and SLO constraints.
@@
Query-Aware Model Serving. Recently, query awareness has shown promises for improving model serving throughput while meeting SLO constraints. In query-aware model serving systems, such as FrugalGPT [11], DiffServe [8], Tabi [43], CascadeBert [25], and CascadeServe [24], each query is processed through a model cascade that is a set of models sequenced from lighter weight ones to heavier weight ones. Each query is routed sequentially through the models in the cascade till a response of acceptable quality is obtained. DiffServe [8] exemplifies such a system tailored to diffusion models. Its cascade uses two stages: every prompt is first processed by a lightweight model, and the generated image is assessed by a discriminator. If the image is deemed low-quality, the prompt is rerouted to a heavyweight model. Easier queries only require the first stage of the cascade, saving resources and latency, while harder queries are routed to the second.
@@
Limitations of Current Solutions. The current query-aware serving systems using model cascades have two key limitations. First, model cascades inherently increase both resource usage and latency as some fraction of the queries need to get routed through multiple models. For instance, in a two-model cascade, all prompts are sent to a lightweight model, which wastes resources for queries that are hard and require heavyweight processing regardless. Second, the models chosen for a cascade may not be optimal for all query workloads. For instance, DiffServe relies on a fixed pair of lightweight and heavyweight models, even though different workloads and latency budgets may favor different combinations.
@@
(i) A hybrid routing architecture. We propose a hybrid architecture (Fig. 1) that combines a router and a discriminator to reduce resource wastage and latency. The architecture begins with a router that computes a query difficulty score using linguistic and semantic features, including token rarity, abstractness, spatial relations, etc. Queries that are predicted to be hard are routed directly to a heavyweight model, skipping the lightweight model(s) in the cascade. Queries that are classified as easy are processed by a lightweight model, whose outputs are then evaluated by an image discriminator for both semantic and aesthetic quality. If the generated images fail to meet the quality requirements, the query is sent to the heavyweight model. This hybrid architecture reduces computational overhead and latency for easy queries and avoids wasteful lightweight model inference on hard queries. We theoretically characterize the conditions for the router and discriminator performance such that the hybrid architecture is guaranteed to outperform the router-only and discriminator-only alternatives in response quality and latency.
(ii) Dynamic choice of models. As diffusion models proliferate, each variant offers different trade-offs between latency and quality. Statically fixing the model in a hybrid architecture across all query workload scenarios is inefficient. Therefore, we design a model serving system that dynamically chooses the models in the hybrid architecture to suit the properties of the current query workload.
@@
(iii) Managing cascade complexity. The plethora of diffusion model variants raises the possibility of complex hybrid architecture consisting of many model variants, discriminators and router components. We show that such complexity is unnecessary and it is sufficient to dynamically choose a simple two-stage model cascade with a single router and discriminator to achieve the benefits of query awareness. This observation greatly reduces the space of architectural choices and enables efficient resource management.
@@
(iv) Reducing the complexity of resource management. Although restricting the design to two-stage cascades simplifies the search space, the system must still decide among many configuration options, including the choice of model pairs, router and discriminator thresholds, model placements, and batch sizes. Moreover, the optimal configuration shifts as the query workload and SLO requirements change. Prior systems such as InferLine [12], INFaaS [34], Proteus [7], and DiffServe [8] do not cover the entire search space because they are not designed to operate over the hybrid architecture. HADIS reduces the combinatorial complexity as follows. We use offline profiling to benchmark all pairwise model combinations under different router and discriminator thresholds and generate a cascade configuration lookup table. The table prunes the search space by identifying Pareto-optimal configurations that offer the best quality-latency trade-offs. At runtime, the system leverages this table to select a configuration that meets the desired SLO constraints. To determine the optimal deployment strategy given a workload distribution and system constraints (e.g., GPU count, latency budget, etc), we formulate the configuration selection and resource allocation as an MILP (mixed integer linear programming) problem, jointly optimizing them to maximize overall serving efficiency and quality.
Figure 1: The hybrid architecture for text-to-image diffusion model serving in HADIS. $\theta$: router threshold; $\tau$: discriminator threshold.
@@
  • We propose a hybrid serving architecture that combines a query router and a discriminator, and theoretically derive conditions under which this hybrid architecture achieves performance benefits.
  • We empirically show that two-model cascades are sufficient for effective model scaling in text-to-image serving, and propose adaptive model cascade selection, dynamically selecting optimal pairwise model cascades based on various workload and latency constraints.
  • We address the resource management problem for serving the hybrid architecture through profiling Pareto-optimal configurations offline and formulating a MILP that jointly optimizes architectural configurations and resource allocation.
@@
2 Background and Rationale for Design
@@
Background for T2I diffusion models. Text-to-image (T2I) diffusion models embody a clear quality-latency trade-off (see Figure 2-left): more compute (e.g., deeper networks or transformers, more denoising steps, etc) typically yields better image quality but longer runtime (and latency) and lesser throughput (i.e., fewer queries processed per second). In our setting on L40S GPUs, the spectrum spans sub-second generation (e.g., SDXL-Lightning at 2 steps $\approx$ 0.5s) to tens of seconds (e.g., SD3.5-Large at 50 steps $\approx$ 27s) for an image of $\times1024$ resolution.
@@
Figure 2: Quality-latency tradeoffs of different diffusion model variants (left) and model cascades (right). Quality is measured by Fréchet Inception Distance [22] (FID, lower is better). The cascades are constructed by the model variants colored in orange in the left panel. The orange dashed lines partition latency into three regimes where each regime has a distinct model cascade that yields the frontier points.
@@
Rationale for two-model cascades. In the context of T2I diffusion models, we show that two-model cascades offer comparable quality-latency trade-offs as model cascades with more model variants. Figure 2-right explores the design space of diffusion-model cascades using four model variants (orange stars in the left panel). For each candidate cascade we enumerate routing policies specified by the fraction of queries routed to the heavyweight model(s), and measure the resulting mean latency (x-axis) and FID (y-axis) on 5k prompts. The right panel plots all feasible cascade configurations in grey by varying model combinations and routing policies. The Pareto frontier [^1] A Pareto frontier $P$ is a set of model cascades such that for every $m\in P$ there is no other model cascade $m^{\prime}\in P$ that has both lower FID and latency than $m$. In that sense, all model cascades in the Pareto frontier represent an optimal tradeoff between the two objectives where one objective can be made better only at the cost of the other. obtained by two-model and three-model cascades are colored in red and blue, respectively. Two empirical facts stand out. First, the red and blue frontiers coincide over the relevant latency range: for any latency budget, a two-model cascade attains essentially the same best FID as a three-model cascade. Second, while the blue frontier can fill in a few intermediate points (i.e., slightly finer granularity), it does not deliver significantly better quality at a given latency.
@@
Figure 3: Limitations of fixed model cascades. In both panels, each point represents a distinct two-model cascade. Comparing between low load (left) and high load (right), the best cascade (circled in red) varies as the system workload and the tolerable SLO violation ratio change.
Rationale for dynamic model choice. Given that our system adopts a two-model cascade, it is natural to ask which pair of models can provide the best system performance. We empirically observe that none of the fixed lightweight/heavyweight model pairs remains optimal across varying query demands.
Figure 2-right shows the best two-model cascade configurations vary with the latency budget. It plots the Pareto frontier of two-model cascade configurations built from four candidates in the left panel. The orange dashed lines partition latency into three regimes. Within each regime, the frontier points arise from a different cascade, <SDXL-Lightning, SD3.5-Turbo> at low latency, <SD3.5-Turbo, SD3.5-Medium> at mid latency, and <SD3.5-Medium, SD3.5-Large> at high latency. We observe that different cascades offer different latencies and response qualities. Hence, when the allowed SLO is tighter, a faster cascade is preferred to satisfy deadlines; when the SLO relaxes, cascades of higher quality can be chosen to increase the response quality.
Figure 3 further shows the best two-model cascades vary with the query loads. It compares several two-model cascades under low demand (left panel) and high demand (right panel) with the SLO held fixed. Each cascade serves queries with the resource allocation algorithm in DiffServe [8]. At low load, all pairs keep SLO violations small. <SD3.5-Turbo, SD3.5-Medium> (red star) attains the best violation rate with competitive FID, while <SD3.5-Medium, SD3.5-Large> trades a slight increase in violations for the best FID (grey diamond). Under high load, as capacity constraints surface, pairs involving slower models (i.e., grey diamond) cannot meet throughput demands and incur large violations, whereas <SDXL-Lightning, SD3.5-Turbo> (blue down-pointing triangle) remains within SLO and delivers the best quality–latency balance for that demand.
@@
Rationale for hybrid architecture. We argue for a hybrid query routing architecture (Fig. 1) by noting that both a router and a discriminator are needed and either one alone is insufficient, as they address complementary failure modes of the model cascade. If we want to solely use a router to evaluate the hardness of queries and route queries based on that, the router should be lightweight, introducing negligible overhead, and also accurate such that most of the queries will not be misclassified, i.e., false “hard” queries can cause unnecessary waste of GPU resources, and false "easy" queries can reduce the response quality. While query hardness has been studied in  [44, 15, 27], estimating query hardness for text-to-image generation with a router is still an open and non-trivial research problem. Building a router that classifies queries with high accuracy is difficult.
@@
A discriminator in the model cascade is usually effective in determining the hardness of a query, which is done in prior work [8, 11]. However, if we use a discriminator alone, all queries will go through the lightweight model regardless of hardness, then be evaluated by the discriminator, which brings in additional overhead, especially when the first model of the cascade requires significant computation. For instance, in Fig. 2, when adopting model pair <SD3.5-Medium, SD3.5-Large>, of which the latency is $\sim$13s and $\sim$27s respectively, the additional overhead for the first model of the cascade can be 33% of the execution latency.
@@
We therefore advocate a hybrid architecture, combining both a router and a discriminator. The rule-based router sends obviously hard queries directly to the heavyweight model, while the rest go to the lightweight model. The discriminator then evaluates the outputs from the lightweight model and catches false “easy” cases for re-routing to heavyweight models. This design avoids unnecessary first-stage work for clearly hard queries, preserves quality by correcting router mistakes, and keeps the added control cost minimal. Our evaluation in §5.4 shows this hybrid routing reduces SLO violations versus discriminator-only and improves FID versus router-only, delivering the best quality-latency trade-off under identical hardware resources.
3 Theoretical Analysis of Hybrid Architecture
This section formalizes the rationale for our hybrid architecture. We theoretically show that there exist parameter settings under which the hybrid architecture (HY) achieves quality better than Router-Only (RO) and Discriminator-Only (DO) at the same expected cost. We then interpret the resulting conditions geometrically as feasibility regions in the ROC planes of the router and the discriminator (see Fig. 4 and Fig. 5).
Preliminary. The analysis considers a hybrid architecture with four components: a cheap, fast model $M_{0}$, a slower, high-quality model $M_{1}$, a router, and a discriminator. The router decides whether each query is sent to $M_{0}$ or directly to $M_{1}$, while the discriminator decides, for queries that went through $M_{0}$, which outputs are good enough to return and which should be escalated to $M_{1}$, as illustrated in Fig. 1. Let $q$ denote the fraction of easy queries in the workload, so hard queries occupy mass 1q1-q. Let $s_{0}$ and $s_{1}$ be the latencies of running $M_{0}$ and $M_{1}$ with $s_{0}<s_{1}$. We define the latency ratio $\alpha=\frac{s_{0}}{s_{1}}\in(0,1)$ which captures how much cheaper the lightweight model is relative to the heavyweight model.
For the router with threshold $\theta$, we define $r_{TP}^{\theta}$ as the probability an easy query is routed to $M_{0}$ (i.e., True Positive Rate, TPR), and $r_{FP}^{\theta}$ as the probability a hard query is routed to $M_{0}$ (i.e., False Positive Rate, FPR). As $\theta$ varies, the pair $(r_{FP}^{\theta},r_{TP}^{\theta})$ traces the router’s ROC curve. Similarly, for the discriminator with threshold $\tau$, we define $d_{TP}^{\tau}$ as the probability that an output from $M_{0}$ on an easy query is accepted (TPR), whereas $d_{FP}^{\tau}$ is the probability that an output on a hard query is incorrectly accepted and completed at $M_{0}$ (FPR). As $\tau$ varies, the pair $(d_{FP}^{\tau},d_{TP}^{\tau})$ traces the discriminator’s ROC curve. Together, these two ROC curves describe the operating space in which we will derive conditions for HY to dominate RO and DO under equal expected cost.
To reason about quality analytically, we define quality using a simple probabilistic model. We assume the heavyweight model $M_{1}$ always produces acceptable quality for both easy and hard queries, while the lightweight model $M_{0}$ is adequate only for easy queries. Under this abstraction, the only way to harm quality is to return an $M_{0}$ output on a hard query. The quality ($Q$) is hence defined as one minus the probability of this failure event. We model cost ($C$) as the expected latency for an architecture to process a query.
Comparison between HY and RO. In a RO architecture, a hard query is mishandled if and only if the router classifies it as easy and sends it to $M_{0}$, which occurs with probability $(1-q)\,r_{FP}^{\theta_{R}}$, yielding:
${\text{Quality:}\quad}Q_{RO}$ $=1-(1-q)r_{FP}^{\theta_{R}},$
${\text{Cost:}\quad}C_{RO}$ $=r_{0}^{\theta_{R}}s_{0}+r_{1}^{\theta_{R}}s_{1},$
$s.t.\quad r_{0}^{\theta_{R}}$ $=qr_{TP}^{\theta_{R}}+(1-q)r_{FP}^{\theta_{R}},\quad r_{1}^{\theta_{R}}=1-r_{0}^{\theta_{R}},$
where $\theta_{R}$ is the router threshold of RO architecture; $r_{0}$ and $r_{1}$ are the fractions of queries routed to $M_{0}$ and $M_{1}$, respectively.
In a HY architecture, a hard query is mishandled only if it is routed to $M_{0}$ and the discriminator incorrectly accepts the $M_{0}$ output. This happens with probability $(1-q)\,r_{FP}^{\theta_{H}}d_{FP}^{\tau_{H}}$:
${\text{Quality:}\quad}Q_{HY}$ $=1-(1-q)r_{FP}^{\theta_{H}}d_{FP}^{\tau_{H}},$
${\text{Cost:}\quad}C_{HY}$ $=r_{0}^{\theta_{H}}s_{0}+(r_{1}^{\theta_{H}}+\pi^{\theta_{H},\tau_{H}})s_{1},$
$s.t.\quad r_{0}^{\theta_{H}}$ $=qr_{TP}^{\theta_{H}}+(1-q)r_{FP}^{\theta_{H}},\quad r_{1}^{\theta_{H}}=1-r_{0}^{\theta_{H}},$
$\pi^{\theta_{H},\tau_{H}}$ $=qr_{TP}^{\theta_{H}}(1-d_{TP}^{\tau_{H}})+(1-q)r_{FP}^{\theta_{H}}(1-d_{FP}^{\tau_{H}}),$
where $\pi$ is the fraction of queries that are first routed to $M_{0}$ and then rerouted to $M_{1}$.
To compare HY against RO, we impose the equal-cost constraint $C_{HY}=C_{RO}$ and analyze the quality gap. Note that only if $\theta_{H}>\theta_{R}$ can the cost be equal. Requiring $Q_{HY}\geq Q_{RO}$ (i.e., HY is at least as good as RO at the same cost) yields the following bounds on the discriminator ROC point $(d_{FP}^{\tau_{H}},d_{TP}^{\tau_{H}})$ of the HY architecture:
$d_{TP}^{\tau_{H}}\geq$ $\frac{\alpha\left(q(r_{TP}^{\theta_{H}}-r_{TP}^{\theta_{R}})+(1-q)(r_{FP}^{\theta_{H}}-r_{FP}^{\theta_{R}})\right)+qr_{TP}^{\theta_{R}}}{qr_{TP}^{\theta_{H}}},$ (1)
$d_{FP}^{\tau_{H}}\leq$ $\frac{r_{FP}^{\theta_{R}}}{r_{FP}^{\theta_{H}}}.$
Figure 4: Feasibility region for the discriminator of HY architecture so that HY outperforms RO. Higher $\alpha$ means smaller gaps between lightweight and heavyweight models.
Figure 5: Feasibility region of the router of HY architecture so that HY outperforms DO. With $\alpha$ increases, the region becomes larger.
The inequalities define a feasibility region in the ROC plane for configuring $\tau_{H}$ : as long as HY employs a discriminator whose $(\text{FPR},\text{TPR})$ lies inside this region, the HY architecture will match RO’s expected cost while achieving better response quality. We visualize such a region in Fig. 4 by fixing $q=0.5$ and traversing the feasible $(r^{\theta}_{TP},r^{\theta}_{FP})$ pairs. As $\alpha$ increases from $0.1$ to $0.2$, the feasible region shrinks. This is because when the lightweight model becomes more expensive relative to $M_{1}$, only discriminator thresholds with sufficiently high $d_{TP}$ and low $d_{FP}$ (i.e., making accurate re-routing decisions) allow HY to still dominate RO at equal cost. In practice, the discriminator ROC lies well above the router ROC, making these conditions practically attainable.
Comparison between HY and DO.. We now compare the hybrid architecture against a DO architecture, where every query first runs through $M_{0}$ and a discriminator with threshold $\tau_{D}$ decides whether to accept the $M_{0}$ output or reroute the query to $M_{1}$.
${\text{Quality:}\quad}Q_{DO}$ $=1-(1-q)d_{FP}^{\tau_{D}},$
${\text{Cost:}\quad}C_{DO}$ $=s_{0}+\pi^{\tau_{D}}s_{1},$
$s.t.\quad\pi^{\tau_{D}}$ $=q(1-d_{TP}^{\tau_{D}})+(1-q)(1-d_{FP}^{\tau_{D}}).$
Requiring $Q_{HY}\geq Q_{DO}$ and $C_{HY}=C_{DO}$ yields bounds on the router ROC point of our hybrid architecture (i.e., $\langle r_{FP}^{\theta_{H}},r_{TP}^{\theta_{H}}\rangle$) for a given discriminator ROC and latency ratio $\alpha$. With $d_{TP}^{\tau_{H}}>\alpha$, we obtain
$r_{TP}^{\theta_{H}}\geq\frac{\alpha(1-q)r_{FP}^{\theta_{H}}+qd_{TP}^{\tau_{D}}-\alpha}{q(d_{TP}^{\tau_{H}}-\alpha)},\quad r_{FP}^{\theta_{H}}\leq\frac{d_{FP}^{\tau_{D}}}{d_{FP}^{\tau_{H}}}$ (2)
The above inequalities define a feasibility region in the router ROC plane: if HY employs a router whose $(\text{FPR},\text{TPR})$ lies inside this region, HY is guaranteed to match DO’s expected cost while achieving better response quality. Fig. 5 illustrates how the required router accuracy depends on $\alpha$. When the lightweight model is very cheap ($\alpha=0.1$), DO can already exploit $M_{0}$ aggressively at low cost, so the feasibility region is relatively small and the router must be quite accurate to provide additional benefit. As $\alpha$ increases, the region expands, indicating that even a moderately accurate router is sufficient for HY to dominate DO.
4 HADIS Design